# Rolling Diffusion Models

### v1

@@ 3.2 A local perspective
3.3 Boundary conditions
While framing rolling diffusion purely from a local perspective is convenient for training and sampling, it introduces complications at the boundaries of the sliding window. That is, given, e.g., the linear local time reparameterization $t^{\text{lin}}_{w}$, we have that given the diffusion time $t$ running from $1$ to 00, the local times run from $(\frac{1}{W},\frac{2}{W},\ldots\frac{W}{W})$ to $(\frac{0}{W},\frac{1}{W},\ldots,\frac{W-1}{W})$. Visually, in Figure 1, the rolling sampling procedure can be seen as moving the sliding window over the diagonal linearly from top left to bottom right such that the local times of the frames remain invariant upon shifting. However, this means that placing the window at the very left edge still results in having partially denoised frames. This means that in this local setting, the signal-to-noise ratios are never minimal, i.e., at full noise.
To account for this, we co-train the rolling diffusion model with an additional schedule that can handle this boundary condition.
$t_{w}^{\text{init}}:=\operatorname{clip}\left(\frac{w}{W}+t\right)$ (23)
This init noise schedule can start from random noise and generates a video in the “rolling state”. Note that this schedule cannot be used as a sliding window noise schedule, however, it contains $t_{w}^{\text{lin}}$ as we will see later. That is, at diffusion time $t=1$, this will put all frames to maximum noise, and at $t=0$ the frames will be in the rolling state. To be precise, it starts from local times $(1,1,\ldots 1)$ and denoises to $(0,\frac{1}{W},\frac{2}{W},\dots,\frac{W-1}{W})$, after which we can start the previously described local rolling diffusion process. From a visual perspective, in Figure 1, this corresponds to placing the window at the upper left corner and moving it down vertically, until it reaches the rolling state that can be used to continue diagonally.
On the domain $[0,\frac{1}{W}]$, this schedule contains the previous local schedule $t_{w}^{\text{lin}}$ as a special case This means that the model could be trained solely with $t_{w}^{\text{init}}$ to handle the boundaries as well as being able to roll out indefinitely. Note, however, that the $t^{\text{lin}}_{w}$ schedule only gets selected $1/W$ of the time during training (assuming $t\sim U(0,1)$). In contrast, this schedule is used almost exclusively at test time, with the exception being the boundary condition on the first $W$ frames. As such, we found it beneficial to include both schedules during training. This blend is achieved by sampling one schedule or the other based on a Bernoulli hyperparameter $\beta$ controlling the probability of selecting between the two schedules.
@@ 3.4 Local training
We consider any ${\bm{x}}\in\mathbb{R}^{D\times W}$, i.e., we are now chunking videos into blocks of $W$ frames From $t_{w}$ one can compute $\alpha_{t_{w}}$ and $\sigma_{t_{w}}$ using any SNR schedule and the aforementioned reparameterizations. Note that when $n_{\text{cln}}>0$ we implicitly get extra clean conditioning. Let ${\bm{z}}\in\mathbb{R}^{D\times W}$.
@@
The objective now becomes.
@@
The training and sampling procedures are summarized in Algorithm 2, Algorithm 3, Algorithm 1. Furthermore, we provide a visual of the rolling sampling loop in Figure 6.
@@
  Require: $p_{\theta}$, $n_{\text{cln}}$, ${\bm{z}}_{0}$ with local diffusion times $(0/W,\dots,(W-1)/W)$, as e.g. given by Algorithm 3.
@@ 4.1 Video diffusion
Methods that specifically treat test-time unrolling of video generation include Yang et al. (2023); Harvey et al. (2022). It was shown that directly parameterizing the conditional distribution of future frames given past frames is preferable (Harvey et al., 2022; Tashiro et al., 2021) as opposed to adapting the denoising schedule of an unconditional diffusion model. Note that these previous approaches never explicitly introduced a notion of time in their training procedure, as opposed to Rolling Diffusion. Specifically, Harvey et al. (2022) compare various such conditioning schemes, but do not explicitly consider a temporally adapted noise schedule.
@@ 4.2 Other time-series diffusion models
Apart from video, sequential diffusion models have also been applied to other time-series data, such as audio (Kong et al., 2021), text (Li et al., 2022), but also scientifically to weather data (Price et al., 2023) or fluid mechanics (Kohl et al., 2023). Lippe et al. (2023) show that incorporating a diffusion-inspired denoising procedure can help recover high frequency information that typically gets lost when using learned numerical PDE solver emulators. Finally, Wu et al. (2023) also study autoregressive models with specialized noising schedules, focusing mostly on text generation.
@@ 5 Experiments
We conduct experiments using data from various domains and explore several conditioning settings. In all our experiments, we use the Simple Diffusion architecture (Hoogeboom et al., 2023) with equal parameters for both standard and rolling diffusion. Specifically, we use two-dimensional spatial convolution blocks after which, in the deepest layers, we have transformer blocks that operate (i.e., attend) both spatially and temporally.
@@ 5.1 Kolmogorov Flow
First, we run an experiment on simulated fluid dynamics from JaxCFD (Kochkov et al., 2021; Dresdner et al., 2022). Specifically, we use the Kolmogorov flow, an instance of the incompressible Navier-Stokes equations. Recently, there has been increasing interest in emulating classical numerical PDE integrators with machine learning models. Various results have shown that these have the capacity to simulate from initial conditions complex systems to high precision (e.g., Li et al. (2020)). To similar ends, generative models are of increasing interest, as they provide several benefits. First, they provide a way to directly obtain marginal distributions over a future state of a physical system, as opposed to numerically rolling out an ensemble of initial conditions. This especially has use-cases in weather or climate modeling, fluid mechanics analyses, and stochastic differential equation studies. Second, they can improve modeling high data frequencies over approaches based on mean-squared error objectives.
The simulation is based on the following partial differential equation $\frac{\partial\mathbf{u}}{\partial\tau}+\nabla\cdot(\mathbf{u}\otimes\mathbf{u})=\nu\nabla^{2}\mathbf{u}-\frac{1}{\rho}\nabla p+\mathbf{f}$, with $\mathbf{u}:[0,\mathcal{T}]\times\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$ is the solution, $\otimes$ the tensor product, $\nu$ the kinematic viscosity, $\rho$ the fluid density, $p$ the pressure field, and, finally, $\mathbf{f}$ the external forcing. We randomly sample viscosities $\nu$ between $5\cdot 10^{-4}$ and $5\cdot 10^{-3}$, and densities between $0.5$ and $2$, making the task of predicting the dynamics non-deterministic. The ground truth data is generated using a finite volume-based direct numerical simulation (DNS) with a maximal time step of $0.05$s with 6 000 time-steps, corresponding to 300 seconds of simulation time, subsampled every 1.5s. Note that this is much longer than typical simulations, which only consider simulation times in the order of tens of seconds. We simulate 200 000 such samples at $64\times 64$, corresponding to 2 terabytes of data. Due to the chaotic nature and the long simulation time, a model cannot be expected to predict the exact future state, which makes it an ideal dataset to test long temporal rollouts. Details on the simulation settings can be found in Appendix B. The model is given 2 input frames containing the horizontal and vertical velocities. Note that the long rollout lengths (up to $\pm$ 60 seconds), together with the large 1.5s strides, make this a more challenging task than the usual “neural emulator” task. E.g., Lippe et al. (2023); Sun et al. (2023) only go up to ±$\pm 15$s with much smaller strides.
@@
To measure how well the generated data matches the ground truth distribution, we propose a method similar to Fréchet Inception Distance (FID) or FVD. We make use of the fact that spatial frequency intensities provide a good summary statistic (Sun et al., 2023; Dresdner et al., 2022; Kochkov et al., 2021). Let ${\bm{f}}_{\bm{x}}\in\mathbb{R}^{F}$ denote a vector of spatial-spectral magnitudes computed using a (two-dimensional) Discrete Fourier Transform (DFT) from a variable ${\bm{x}}$. Then, ${\mathbf{F}}_{\mathcal{D}}\in N\times F$ denotes all the frequencies of a (test-time) dataset $\mathcal{D}:=\{{\bm{x}}_{1},\dots,{\bm{x}}_{N}\}$. Let ${\mathbf{F}}_{\theta}$ denote the Fourier magnitudes of $N$ samples from a generative model. We now compute the Frećhet distance between the generated samples and the true data by setting
@@
In Figure 3 we present FSD computed from the horizontal velocity fuilds of the fluid. Regarding rolling diffusion, we use the $t^{\text{init}}_{w}(n_{\text{cln}})$ reparameterization with $n_{\text{cln}}=2$, and use $t^{\text{lin}}_{w}(n_{\text{cln}})$ for long rollouts. It is clear that an autoregressive MSE-based model, as typically used in the literature, is not suitable for this task. For standard diffusion, we iteratively generate $W-n_{\text{cln}}$ frames, after which we concatenate these to the conditioning and continue the rollout. Rolling diffusion always shifts the window by one, sampling using the process described before. We see that the rolling diffusion is able to consistently outperform the standard diffusion methods, regardless of various conditioning settings and window sizes, denoted with ‘$(n_{\text{cln}},W-n_{\text{cln}})$’. We provide an example rollout in Figure 2, where we plot the vorticity of the velocity field. We numerically present our results in Appendix C.
@@ 5.2 BAIR Robot Pushing Dataset
The results are shown in Table 1. We observe that both standard diffusion and rolling diffusion using the same (Simple Diffusion) architecture outperform previous methods. Additionally, we see that there is no significant difference between the standard and rolling framework in this setting. This is because the sampled sequences are, in both cases, indistinguishable from the true data Figure 4. One could say that the models are can completely solve this task, yielding no significant difference in performance. However, it is still interesting that both models are able to outperform existing methods.
@@
MAGVIT (Yu et al., 2023a) ${}^{\dagger}$
MAGVITv2 (Yu et al., 2023b) ${}^{\dagger}$
W.A.L.T.-L (Gupta et al., 2023) ${}^{\dagger}$
@@
5.3 Kinetics-600
Figure 5: Top: Rolling Diffusion rollout on the Kinetics-600 dataset. Bottom: ground-truth.
Baseline
We compare against previous methods using 5 input frames and 11 output frames, and show the results in Table 2 The evaluation metric is again FVD. We note that many of the current SOTA methods are two-stage, meaning that they use an autoencoder and run diffusion in latent space. While empirically compelling (Rombach et al., 2022), this makes it hard to isolate the effect of the diffusion model itself. Note that it is not always clear whether the autoencoder parameters are included in the parameter count for two-stage methods, or on what data they are pretrained. Running diffusion using the standard diffusion U-ViT architecture achieves an FVD of 3.9, which is comparable to the best two-stage methods. Rolling diffusion has a strong disadvantage in this case: (1) the baseline generates all frames at once; and (2) with a stride of 1, there is very little dynamics in the 16 frames, mostly suitable for a standard diffusion model. Still, rolling diffusion achieves a FVD of 5.2.
@@ 5.3 Kinetics-600
Next, we compare the models’ capabilities to autoregressively rollout and show the results in Table 3. All settings use a window size of $W=16$ frames, where set $n_{\text{cln}}$ to various settings, denoted with ‘cond-gen’. Furthermore, we train rolling diffusion with a mix of the default schedule and a rescaled schedule. The default rolling diffusion schedule $t^{\text{lin}}_{w}$ is oversampled at a rate of $\beta$.
We analyze two settings, one with a stride (also known as frame-skip or frame-step) of 1, rolling out for 64 steps, and another setting with a stride of 8 rolling out for 24 steps, effectively predicting ahead up to the $192$th frame. In the first setting, standard diffusion performs better, quite possibly due to the invariability of the data. We oversample the linear rolling schedule with a rate of $\beta=0.9$ to account for the high number of test-time steps. Rolling diffusion consistently wins in the second setting, which is much more dynamic. Note also that single-frame diffusion significantly underperforms here and that larger block autoregression is favorable. See an example rollout in Figure 5. From Appendix F, we get an indication of the effect of the oversampling rate on the performance of rolling diffusion. In this case, slightly oversampling $t^{\text{lin}}_{w}$ yields the best result.
@@ 6 Conclusion
We presented Rolling Diffusion Models, a new DDPM framework that progressively noises (and denoises) data through time. Validating our method on video and fluid mechanics data, we observed that rolling diffusion’s natural inductive bias gets most effectively exploited when the data is highly dynamic. This allows for exciting future directions in, e.g., video, audio, and weather or climate modeling.
Broader Impact
@@
Appendix A Hyperparameters
In this section we denote the hyperparameters for the different experiments. Throughout the experiments we use U-ViTs which are essentially U-Nets with MLP Blocks instead of convolutional layers when self-attention is used in a block. In the PDE experiments are relatively small architecture is used. For BAIR we used a larger architecture, increasing both the channel count and the number of blocks. For K600 we used even larger architectures, because this dataset turned out to be the most difficult to fit. For all the specifications see Table 4, 5 and 6.

### v2

@@ 3.4 Local training
We briefly discuss training details under the aforementioned local time reparameterizations. We now consider ${\bm{x}}\in\mathbb{R}^{D\times W}$, chunking videos into blocks of $W$ frames. From $t_{w}$ one can compute $\alpha_{t_{w}}$ and $\sigma_{t_{w}}$ using typical SNR schedules. Let ${\bm{z}}\in\mathbb{R}^{D\times W}$, we have as the forward (noising) process
@@
where ${\bm{x}}^{w}$ denotes the $w$-th frame of ${\bm{x}}$.
The training objective becomes
@@
Algorithm 1 Rolling Diffusion: Training
  repeat
    Sample ${\bm{x}}$ from $\mathcal{D}_{\mathrm{tr}}$, $t\sim U(0,1)$, $y\sim B(\beta)$
    if $y$ then
       Compute local time $t_{w}^{\text{init}}(n_{\text{cln}})$, $w=0,\dots,W-1$
    else
       Compute local time $t_{w}^{\text{lin}}(n_{\text{cln}})$, $w=0,\dots,W-1$
    end if
    Compute $\alpha_{t_{w}}$ and $\sigma_{t_{w}}$ for all $w=0,\dots,W-1$
    Compute $\hat{{\bm{x}}}\leftarrow f_{\theta}({\bm{z}}_{t,{\bm{{\epsilon}}}};t)$
    Update $\theta$ using $L_{\text{loc},\theta}({\bm{x}};t,{\bm{{\epsilon}}})$
  until Converged
@@
  Require: $p_{\theta}$, $n_{\text{cln}}$, ${\bm{z}}_{0}$ with local diffusion times $(0/W,\dots,(W-1)/W)$ (i.e., progressively noised).
@@
The training and sampling procedures are summarized in Algorithm 1 and Algorithm 2. We summarize sampling at the boundary using $t^{\text{init}}_{w}$ in Appendix B. Furthermore, we provide a visual of the rolling sampling loop in Figure 1.
@@ 4.1 Video diffusion
Methods that explore autoregressive video generation include Yang et al. (2023); Harvey et al. (2022). Directly parameterizing the conditional distribution of future frames given past frames is preferable (Harvey et al., 2022; Tashiro et al., 2021) compared to adapting the denoising schedule of an unconditional diffusion model. Unlike previous approaches, Rolling Diffusion explicitly introduces a notion of time in the training procedure. Harvey et al. (2022) compare various conditioning schemes but do not explicitly consider a temporally adapted noise schedule.
@@ 4.2 Other time-series diffusion models
Apart from video, sequential diffusion models have also been applied to other modes of time-series data, such as audio (Kong et al., 2021), text (Li et al., 2022), but also scientifically to weather data (Price et al., 2023) or fluid mechanics (Kohl et al., 2023). Lippe et al. (2023) show that incorporating a diffusion-inspired denoising procedure can help recover high frequency information that typically gets lost when using learned numerical PDE solver emulators. Dyffusion (Cachay et al., 2023) employs a forecasting and interpolation model in a two-stage fashion. Both models are trained with MSE objectives. This means that the generative model can be interpreted as factorized Gaussians, potentially leading to blurry predictions when generating stochastic or highly chaotic data as considered in this work. Wu et al. (2023) also study autoregressive models with specialized noising schedules, focusing mostly on text generation. Finally, Zhang et al. (2023) published around the same time as the current work a paper with similar ideas as rolling diffusion. There are, however, some core differences. This work analyzes in more detail how rolling diffusion relates to a global, well-defined diffusion process, motivating why training on a sliding window is acceptable. We isolate the effect of rolling diffusion and study it in more detail, while Zhang et al. (2023) combine local noise levels with additional losses, potentially blurring the effect of the sliding window idea with the impact of these auxiliary terms. Finally, we introduce rolling diffusion schedules that deal with the boundary situations, allowing for generating sequences in an end-to-end manner.
@@ 5 Experiments
We conduct experiments using data from various domains and explore several conditioning settings. In all our experiments, we use the Simple Diffusion architecture (Hoogeboom et al., 2023) with equal parameters for both standard and rolling diffusion. We use two-dimensional spatial convolution blocks after which we have transformer blocks that attend both spatially and temporally in the deepest layers. Hyperparameter settings can be found in Appendix C. A note on the runtime complexity can be found in Appendix D.
@@ 5.1 Kolmogorov Flow
First, we run an experiment on simulated fluid dynamics from JaxCFD (Kochkov et al., 2021; Dresdner et al., 2022). Specifically, we use the Kolmogorov flow, an instance of the incompressible Navier-Stokes equations. Recently, there has been increasing interest in emulating classical numerical PDE integrators with machine learning models. Various results have shown that these have the capacity to simulate from initial conditions complex systems to high precision (e.g., Li et al. (2020)). To similar ends, generative models are of increasing interest, as they provide several benefits. First, they provide a way to directly obtain marginal distributions over a future state of a physical system, as opposed to numerically rolling out an ensemble of initial conditions. This especially has use-cases in weather or climate modeling, fluid mechanics analyses, and stochastic differential equation studies. Second, they can improve modeling high frequency information over approaches based on mean-squared error objectives.
The simulation is based on the following partial differential equation $\frac{\partial\mathbf{u}}{\partial\tau}+\nabla\cdot(\mathbf{u}\otimes\mathbf{u})=\nu\nabla^{2}\mathbf{u}-\frac{1}{\rho}\nabla p+\mathbf{f}$, with $\mathbf{u}:[0,\mathcal{T}]\times\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$ is the solution, $\otimes$ the tensor product, $\nu$ the kinematic viscosity, $\rho$ the fluid density, $p$ the pressure field, and, finally, $\mathbf{f}$ the external forcing. We randomly sample viscosities $\nu$ between $5\cdot 10^{-4}$ and $5\cdot 10^{-3}$, and densities between $0.5$ and $2$, making the task of predicting the dynamics non-deterministic. The ground truth data is generated using a finite volume-based direct numerical simulation (DNS) with a maximal time step of $0.05$s with 6 000 time-steps, corresponding to 300 seconds of simulation time, subsampled every 1.5s. Note that this is much longer than typical simulations, which only consider simulation times in the order of tens of seconds. We simulate 200 000 such samples at $64\times 64$, corresponding to 2 terabytes of data. Due to the chaotic nature and the long simulation time, a model cannot be expected to predict the exact future state, which makes it an ideal dataset to test long temporal rollouts. Details on the simulation settings can be found in Appendix E. The model is given 2 input frames containing the horizontal and vertical velocities. Note that the long rollout lengths (up to $\pm$ 60 seconds), together with the large 1.5s strides, make this a more challenging task than the usual ‘neural emulator’ task, where the predictions can be extremely accurate due to the shorter time-scales and higher temporal resolutions. For example, Lippe et al. (2023); Sun et al. (2023) only go up to ±$\pm 15$s with much smaller strides.
@@
Typical PDE-emulation tasks directly compare models to ground truth using data-space RMSE scores. However, in our uncertain setting we are more interested in how well the generated distribution matches the target distribution. The ground-truth simulator provides no inherent uncertainty estimation, and running ensembles from various initial parameters is not straightforward due to a lack of proposal distribution for these initial settings. Instead, we propose a method similar to Fréchet Inception Distance (FID) or FVD. We make use of the fact that spatial frequency intensities provide a good summary statistic (Sun et al., 2023; Dresdner et al., 2022; Kochkov et al., 2021). Let ${\bm{f}}_{\bm{x}}\in\mathbb{R}^{F}$ denote a vector of spatial-spectral magnitudes computed using a (two-dimensional) Discrete Fourier Transform (DFT) from a variable ${\bm{x}}$. Then, ${\mathbf{F}}_{\mathcal{D}}\in N\times F$ denotes all the frequencies of a (test-time) dataset $\mathcal{D}:=\{{\bm{x}}_{1},\dots,{\bm{x}}_{N}\}$. Let ${\mathbf{F}}_{\theta}$ denote the Fourier magnitudes of $N$ samples from a generative model. We now compute the Frećhet distance between the generated samples and the true data by setting
@@
We provide an example rollout in Figure 3, where we plot the vorticity of the velocity field. Quantitatively, we present in Figure 4 the FSD results derived from the horizontal velocity fields of the fluid. Note that the standard diffusion ($n_{\text{cln}}$-1) baselines can be seen as a adaptations of PDE Refiner (Lippe et al., 2023) and (Kohl et al., 2023) for our task. Regarding rolling diffusion, we use the $t^{\text{init}}_{w}(n_{\text{cln}})$ reparameterization with $n_{\text{cln}}=2$, and use $t^{\text{lin}}_{w}(n_{\text{cln}})$ for long rollouts. It is clear that an autoregressive MSE-based model, as typically used in the literature, is not suitable for this task. For standard diffusion, we iteratively generate $W-n_{\text{cln}}$ frames, after which we concatenate these to the conditioning and continue the rollout. Rolling diffusion always shifts the window by one, sampling using the process described before. Rolling diffusion consistently outperforms standard diffusion methods, regardless of conditioning settings and window sizes $(n_{\text{cln}},W-n_{\text{cln}})$. Additional qualitive and numerical results can be found in Appendix F.
@@ 5.2 BAIR Robot Pushing Dataset
The results are shown in Table 1. We observe that both standard diffusion and rolling diffusion using the same (Simple Diffusion) architecture outperform previous methods. Additionally, we see that there is no significant difference between the standard and rolling framework in this setting. This is because the sampled sequences are, in both cases, indistinguishable from the true data Figure 5.
5.3 Kinetics-600
Figure 6: Top: Rolling Diffusion rollout on the Kinetics-600 dataset. Bottom: ground-truth. License: CC BY 4.0.
Baseline
@@ 5.3 Kinetics-600
MAGVIT (Yu et al., 2023a) 9.9
MAGVITv2 (Yu et al., 2023b) 4.3
W.A.L.T.-L (Gupta et al., 2023) 3.3
@@
We compare against previous methods using 5 input frames and 11 output frames, and show the results in Table 2 The evaluation metric is again FVD. We note that many of the current SOTA methods are two-stage, meaning that they use an autoencoder and run diffusion in latent space. While empirically compelling (Rombach et al., 2022), this makes it hard to isolate the effect of the diffusion model itself. Note that it is not always clear whether the autoencoder parameters are included in the parameter count for two-stage methods, or on what data they are pretrained. Running diffusion using the standard diffusion U-ViT architecture achieves an FVD of 3.9, which is comparable to the best two-stage methods. Rolling diffusion has a strong disadvantage in this case: (1) the baseline generates all frames at once, thereby not suffering from autoregressive errors; and (2) with a stride of 1, there is very little dynamics in the 16 frames, mostly suitable for a standard diffusion model. Still, rolling diffusion achieves a competitive FVD of 5.2.
@@
Next, we compare the models’ capabilities to autoregressively rollout and show the results in Table 3. All settings use a window size of $W=16$ frames, exploring several $n_{\text{cln}}$ settings, denoted with ‘Cond-Gen’. During training, we mix $t^{\text{lin}}_{w}$ with a slightly adjusted version of $t^{\text{lin}}_{w}$ (see Appendix G) at an oversampling rate of $\beta$.
We analyze two settings, one with a stride (also known as frame-skip or frame-step) of 1, rolling out for 64 steps, and another setting with a stride of 8 rolling out for 24 steps, effectively predicting ahead up to the $192$th frame. In the first setting, standard diffusion performs better, quite possibly due to the invariability of the data. We oversample the linear rolling schedule with a rate of $\beta=0.9$ to account for the high number of test-time steps. Rolling diffusion consistently wins in the second setting, which is much more dynamic. Note also that single-frame diffusion significantly underperforms here and that larger block autoregression is favorable. See an example rollout in Figure 6. Additionally, we compare to TECO (Yan et al., 2023), which uses slightly different input-output settings, in Appendix F. From Appendix H, we get an indication of the effect of the oversampling rate on the performance of rolling diffusion. In this case, slightly oversampling $t^{\text{lin}}_{w}$ yields the best result.
From these (and previous) results we draw the conclusion that rolling diffusion is particularly effective in dynamic settings, where the data is highly variable.
@@ 6 Conclusion
We presented Rolling Diffusion Models, a new DDPM framework that progressively noises (and denoises) data through time. Validating our method on video and fluid mechanics data, we observed that rolling diffusion’s natural inductive bias gets most effectively exploited when the data is highly dynamic. In this setting, Rolling Diffusion outperforms various existing methods. This allows for exciting future directions in, e.g., video, audio, and weather or climate modeling.
Impact Statement
@@
Appendix A Rolling Diffusion Objective
Standard Diffusion
Further, when $T\rightarrow\infty$ we get continuous analog of Equation 30 (Kingma et al., 2021). In practice, instead of using the resulting KL objective, one typically uses a weighted loss, which in some cases still corresponds to an importance-weighted KL (Kingma & Gao, 2023).
$\mathcal{L}_{w}$ $:=\frac{1}{2}\mathbb{E}_{t\sim U(0,1),{\bm{{\epsilon}}}\sim\mathcal{N}(0,1)}\left[w(\lambda_{t})\cdot-\frac{d\lambda_{t}}{dt}\lVert\hat{{\bm{{\epsilon}}}}_{\theta}({\bm{z}}_{t};\lambda_{t})-{\bm{{\epsilon}}}\rVert^{2}\right]\,,$ (31)
where $\lambda_{t}:=\log\operatorname{SNR}(t)$. Note that in the main paper we directly write the weighting factor expression as $a(t)$ for ease of notation. The weighting function is often changed to improve image quality, for instance by being $-1/\frac{d\lambda_{t}}{dt}$ so that the objective becomes the simple $\epsilon$-loss.
Rolling Diffusion Objective
In rolling diffusion, the signal-to-noise schedule is unchanged but one has to account for the local time reparameterization. Recall that $t_{k}:=t_{k}(t)$ denotes the local time reparameterization. Using similar derivations as (Kingma et al., 2021), where we can factorize the KL divergence also over frame indices, we get the rolling diffusion loss:
$\mathcal{L}_{\infty}$ $=\mathbb{E}_{t\sim U(0,1),{\bm{{\epsilon}}}\sim\mathcal{N}(0,1)}\left[\sum_{k=1}^{K}w(\lambda(t_{k}))\cdot-\frac{d\lambda(t_{k})}{dt_{k}}\frac{dt_{k}}{dt}\cdot\lVert{\bm{{\epsilon}}}^{k}-\hat{{\bm{{\epsilon}}}}^{k}_{\theta}({\bm{z}}_{{\bm{{\epsilon}}},t};t)\rVert^{2}\right]\,,$ (32)
where we again can apply a custom weighting factor.
Recall the frame categorization of the main paper, i.e.,
$\text{clean}(s,t)$ :=$:=\{k\mid s_{k}=t_{k}=0\}\,,$ (33)
$\text{noise}(s,t)$ :=$:=\{k\mid s_{k}=t_{k}=1\}\,,$ (34)
$\text{win}(s,t)$ :=$:=\{k\mid s_{k}\in[0,1),t_{k}\in(s_{k},1]\}\,,$ (35)
$\mathcal{L}_{\infty}$ $=\mathbb{E}_{t\sim U(0,1),{\bm{{\epsilon}}}\sim\mathcal{N}(0,1)}\left[\sum_{k\in\text{win}(t)}w(\lambda(t_{k}))\cdot-\frac{d\lambda(t_{k})}{dt_{k}}\frac{dt_{k}}{dt}\cdot\lVert{\bm{{\epsilon}}}^{k}-\hat{{\bm{{\epsilon}}}}^{k}_{\theta}({\bm{z}}_{{\bm{{\epsilon}}},t};t)\rVert^{2}\right]\,.$ (36)
Since the weighting function $w$ is arbitrary, we can choose it such that the entire prefactor vanishes, resulting in the typical $\epsilon$-loss. However, the variational bound tells us that we should pay no price for any error made on the noise or clean frames. Again, in the main paper we replace all the prefactors with $a(t_{k})$ for notational convenience.
Appendix B Algorithms
Algorithm 3 Rolling Diffusion: sampling at the boundary.
  ${\bm{z}}_{1}\leftarrow\operatorname{concat}({\bm{x}},{\bm{z}}_{1})$
  for $t=1,(T-1)/T,\dots,1/T$ do
    Compute local times $t_{w}^{\text{init}}(n_{\text{cln}})$, $w=0,\dots,W-1$
    Sample ${\bm{z}}_{t-1/T}\sim p_{\theta}({\bm{z}}_{t-1/T}|{\bm{z}}_{t})$
  end for
  Return ${\bm{z}}_{0}$, which now has local times $(0/W,1/W,\dots,(W-1)/W)$
Appendix C Hyperparameters
In this section we denote the hyperparameters for the different experiments. Throughout the experiments we use U-ViTs which are essentially U-Nets with MLP Blocks instead of convolutional layers when self-attention is used in a block. In the PDE experiments relatively small architectures are used. For BAIR, we used a larger architecture, increasing both the channel count and the number of blocks. For K600 we used even larger architectures, because this dataset turned out to be the most difficult to fit.